Calculation of individual costs and WACC - Dillon Labs has asked its financial manager to measure the cost of each specific type of capital as well as the common stock equity (retained earnings, new common stock, or both). The firm's tax rate is 29%.

Debt - the firm can sell for $1010 a 15-year, $1000-par value bond paying annual interest at a 6.00% coupon rate. A flotation cost of 3.5% of the par value is required.

Preferred stock - 9.50% (annual dividend) preferred stock having a par value of $100 can be sold for $98. An additional fee of $6 per share must be paid to the underwriters.

Common stock - the firm's common stock is currently selling for $59.43 per share. The stock has paid a dividend that has gradually increased for many years, rising from $2.70 ten years ago to the $4.40 dividend payment D0, that the company just recently made. If the company wants to issue new new common stock, it will sell them $1.50 blow the current market price to attract investors, and the company will pay $3.50 per share in flotation costs.

a. calculate the after-tax cost of debt.

b. calculate the cost of preferred stock.

c. calculate the cost of common stock (both retained earnings and new common stock).

d. calculate the WACC for Dillon Labs.
Answer #1

a)
FV = 1000
PV = 1010 * (1 - 3.5%) = 974.65
Nper = 15
PMT = 1000 * 6% = 60

Before tax cost of debt can be calculated by using the following excel formula:
=RATE(nper,pmt,pv,fv)
=RATE(15,60,-974.65,1000)
= 6.27%

After tax cost of debt = Before tax cost of debt * (1 - tax rate)
= 6.27% * (1 - 0.29)
= 4.45%

After tax cost of debt = 4.45%

b)
Annual dividend = 100 * 9.50% = $9.50

Cost of preferred stock = Annual dividend / (Share value - floatation cost)
= $9.50 / ($98 - $6)
= $9.50 / 92
= 10.33%

c)
Dividend year 1 = $2.70
Dividend year 10 = $4.40

Growth rate can be calculated by using the following excel formula:
=RATE(nper,pmt,pv,fv)
=RATE(10,0,-2.70,4.40)
= 5%

Next year dividend = $4.40 * (1 + 5.%) = $4.62

Cost of common stock = Next year dividend / (Share value - costs) + growth rate
= $4.62 / ($59.43 - $1.50 - $3.50) + 5%
= $4.62 / $54.43 + 5%
= 13.49%

Cost of retained earnings = Next year dividend / Share value + growth rate
= $4.62 / $59.43 + 5%
= 12.78%

Average cost of equity = (13.49% + 12.78%) / 2 = 13.14%